The paint used to make lines on roads must reflect enough light to be clearly visible at night. Let μ denote the true average reflectometer reading for a new type of paint under consideration A test of Ho: μ 20 versus Ha: μ > 20 will be based on a random sample of size n from a normal population distribution. What conclusion is appropriate in each of the following situations? (Round your P-values to three decimal places.) (a) z3.3, a-0.05 P-value - State the conclusion in the p roblem context. O Reject the null hypothesis. There is sufficient evidence to conclude that the new paint has a reflectometer reading higher than 20 O Do not reject the null hypothesis. There is sufficient evidence to conclude that the new paint has a reflectometer reading higher than 20 O Reject the null hypothesis. There is not sufficient evidence to conclude that the new paint has a reflectometer reading higher than 20 O Do not reject the null hypothesis. There is not sufficient evidence to conclude that the new paint has a reflectometer reading higher than 20 (b) z-1.8, α# 0.01 P-value- State the conclusion in the problem context. O Reject the null hypothesis. There is not sufficient evidence to conclude that the new paint has a reflectometer reading higher than 20 O Do not reject the null hypothesis. There is sufficient evidence to conclude that the new paint has a reflectometer reading higher than 20 O Reject the null hypothesis. There is sufficient evidence to conclude that the new paint has a reflectometer reading higher than 20 O Do not reject the null hypothesis. There is not sufficient evidence to conclude that the new paint has a reflectometer reading higher than 20 (c) z-_0.3 , α-0.10 P-value- State the conclusion in the problem context. O Do not reject the null hypothesis. There is sufficient evidence to conclude that the new paint has a reflectometer reading higher than 20 O Reject the null hypothesis. There is not sufficient evidence to conclude that the new paint has a reflectometer reading higher than 20 O Do not reject the null hypothesis. There is not sufficient evidence to conclude that the new paint has a reflectometer reading higher than 20 O Reject the null hypothesis. There is sufficient evidence to conclude that the new paint has a reflectometer reading higher than 20
Answer #1

Given that, the null and the alternative hypotheses are,

Ho : μ = 20 versus Ha : μ > 20

a) z = 3.3, a-0.05

p-value = P(Z > 3.3) = 1 - P(Z < 3.3) = 1 - 0.9995 = 0.0005

p-value = 0.001 ( rounded to three decimal places)

Here, p-value = 0.001 < a-0.05

so, we reject H0.

Conclusion: Reject the null hypothesis. There is sufficient evidence to conclude that the new paint has a reflectometer reading higher than 20.

b) z = 1.8, α 0.01

p-value = P(Z > 1.8) = 1 - P(Z < 1.8) = 1 - 0.9641 = 0.0359

p-value = 0.036 ( rounded to three decimal places)

Here, p-value = 0.036 > α 0.01

so, we fail to reject H0.

Conclusion: Do not reject the null hypothesis. There is not sufficient evidence to conclude that the new paint has a reflectometer reading higher than 20.

c) z = -0.3, α = 0.10

p-value = P(Z > -0.3) = P(Z < 0.3) = 0.6179

p-value = 0.618 ( rounded to three decimal places)

Here, p-value = 0.618  > α = 0.10

so, we fail to reject H0.

Conclusion: Do not reject the null hypothesis. There is not sufficient evidence to conclude that the new paint has a reflectometer reading higher than 20.